Jordan is 12 years older than Anna the sum of their ages is 50 find the age of each person
Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

Jordan is 12 years older than Anna the sum of their ages is 50 find the age of each person
================
J - A = 12
J + A = 50
-------------------------- Add

Answer by greenestamps(13334)



Note if a formal algebraic solution is not required, there is a common strategy that can make problems like this easy to solve mentally.

Think of rewording the problem to say the ages differ by 12 years and their sum is 50. Then reason as follows:

(1) Two ages that are the same and have a sum of 50 are 25 and 25.
(2) To get two ages that differ by 12, take the two 25s and add 6 to one and subtract 6 from the other.

That gives you the answer: the ages are 25+6=31 and 25-6=19.